x^4 - 10x^2 + 9 = 0
Answer by JulietG(1812)

x^4 - 10x^2 + 9 = 0
To factor, we know that the first factor will be x^2
(x^2 )(x^2 )
The second number factors to 9 -- so it's either 9,1 or 3,3. 3 +/- 3 is 6 or 0, so that's not it. Therefore, it's 9 and 1.
(x^2 9)(x^2 1)
To find the correct sign, look at the middle number. What combination of +9, -9 and +1, -1 will equal -10 when added?
(x^2-9)(x^2-1)
Put them to zero to find the possible X solutions.
x^2-9=0
Add 9 to each side
x^2=9
x=3 or -3 (when you square it, either works)
The second solution
x^2-1=0
Add 1 to each side
x^2=1
x= 1 or -1